# Correlated interlayer quantum Hall state in alternating twisted trilayer graphene

### v1

@@ VI Discussion
A further characteristic of this system is the emergence of an interlayer excitonic quantum Hall state at $\nu_{\rm tot}=-1$. In this regime the top sheet forms a conventional integer quantum Hall state at $\nu_{\rm top}=-2$, while the middle and bottom sheets are simultaneously half filled, creating conditions for exciton condensation across adjacent layers. The resulting suppression of $R_{xx}$ mirrors excitonic states reported previously in twisted bilayer and double bilayer graphene, but here it coexists with a fully filled outer layer, providing a new trilayer realization of a bilayer hallmark. The field dependence of this feature confirms its interaction-driven origin. In summary, ATTG combines conventional quantum Hall features with emergent interlayer correlated phases, highlighting correlated interlayer quantum Hall phases unique to trilayer geometries with tunable layer asymmetry.
VII Methods
Device fabrication
We fabricated a graphite/h-BN/twisted trilayer graphene/h-BN/graphite/h-BN (from bottom to top) using dry pick up method with Elvacite stamp. Both top and bottom graphite were used as gates. The assembled heterostructure was patterned into a Hall bar geometry using electron-beam lithography followed by reactive ion etching with mixture of $\text{CF}_{4}$ and $\text{O}_{2}$ at power of 40 W. Finally, a metal electrode was deposited with Cr/Au (5/100nm) by e-beam evaporator.
Transport measurements
Transport measurements were performed with the help of low frequency lock-in techniques using an ac current of 100 nA and a frequency of 13.333 Hz in a He4 refrigerator with a base temperature of 1.5 K. In order to reduce contact resistance in areas not covered by the graphite gates, the silicon substrate was biased at $\pm$70 V during the measurements.
Capacitance model
The total filling factor $\nu_{\text{tot}}$ and the displacement field $D/\epsilon_{0}$ were determined within a simple capacitance model as
$\nu_{\text{tot}}=\frac{n_{\text{tot}}h}{eB}=\frac{(C_{\text{t}}V_{\text{t}}+C_{\text{b}}V_{\text{b}})h}{e^{2}B},\qquad\frac{D}{\epsilon_{0}}=\frac{C_{\text{t}}V_{\text{t}}-C_{\text{b}}V_{\text{b}}}{2\epsilon_{0}}.$
Here $n_{\text{tot}}$ is the carrier density, $C_{\text{t(b)}}$ the capacitance per unit area of the h-BN dielectric between the top (bottom) gate and the trilayer graphene, and $V_{\text{t(b)}}$ the applied top (bottom) gate voltage [33, 35, 38, 39, 41]. $e$ denotes the elementary charge, $h$ Planck’s constant, and $B$ the magnetic field.
Tight-binding calculation
Following our approach for magic-angle alternating twist graphenes [42], we here too model the system within a tight-binding framework in the localized orbital basis $|i\rangle$} of carbon sites [43]:
$\hat{H}=\sum_{i}\epsilon_{i}|i\rangle\langle i|+\sum_{i,j}t_{ij}|i\rangle\langle j|,$ (1)
where $\epsilon_{i}$ are onsite terms from the F2G2 graphene parametrization [44], corrected by the onsite energy shifts $\Delta$ and $\delta$ as discussed in the main text, and $t_{ij}$ are distance-dependent hopping amplitudes. In our SHE model [45], intralayer couplings follow the F2G2 model, while interlayer tunneling is described by the two-center approximation:
$t_{ij}^{\rm inter}=S,\exp\left(\frac{c_{ij}-p}{q}\right)t^{\rm TC}_{ij},$ (2)
with p=3.25p=3.25Å, q=1.34q=1.34Å, and $c_{ij}$ the vertical separation. The two-center term is
$t^{\rm TC}{ij}=V{pp\pi}(r_{ij})\left[1-\left(\frac{c_{ij}}{r_{ij}}\right)^{2}\right]+V_{pp\sigma}(r_{ij})\left(\frac{c_{ij}}{r_{ij}}\right)^{2},$ (3)
where $r_{ij}=|{\bm{r}}_{ij}|$, with
$V_{pp\pi}(r)=V_{pp\pi}^{0}\exp\left(-\frac{r-a_{0}}{r_{0}}\right)$ (4)
and
$V_{pp\sigma}(r)=V_{pp\sigma}^{0}\exp\left(-\frac{r-c_{0}}{r_{0}}\right).$ (5)
We set $a_{0}=1.42$Å, $c_{0}=3.35$Å, $V_{pp\pi}^{0}=-2.7$eV, $V_{pp\sigma}^{0}=0.48$eV, and $r_{0}=0.184a$. Interactions vanish beyond 4.94.9Å in accordance with Ref. [46]. The prefactor $S$ is commonly adjusted to reproduce the first magic-angle near $1.1^{\circ}$[47]. Since our focus is on a large angle of $\sim 5^{\circ}$, where moiré reconstruction is negligible, we adopt the rigid-lattice value $S=0.895$ without further correction. For completeness, we note that energy minimization relaxations, using LAMMPS [48] with REBO2 intralayer[49] and DRIP interlayer [50], fitted to EXX-RPA data [51], produce only minor modifications at such large twist and do not affect the observables considered. For electronic properties, we apply the Lanczos method with random-phase vectors to compute the density of states [52]:
[…]
with broadening parameter $\eta$. Magnetic fields are included via a Peierls substitution of phases in the hopping amplitudes [53].
Mean-field Hartree-Fock calculation
We theoretically model our system of ATTG with the continuum Dirac Hamiltonian consisting of the phenomenological short-range interaction terms [35], potential energy terms induced by the displacement fields and Zeeman effect, and the capacitive energy term modeled by three infinite plate (See Supplementary Note 4 of SI for details). We focus on the 12-dimensional lowest Landau levels, given by layer $\times$ valley $\times$ spin = $3\times 2\times 2$. We determine the ground state using the Hartree–Fock method within a valley-, layer-, and spin-polarized ansatz. Applying Hartree-Fock method and Fierz identity together, we reduce the interactions into a sum of one-body pseudospin operators expressed in the direct product of Pauli-matrix basis, and represent the state by 12-component spinor [54]. We performed a brute-force search restricted to diagonal projectors. Further details are provided in Supplementary Note 4 of the SI.
VIII Data availability
Source data for all main text figures are included in the Supplementary Information. Any other data supporting the findings of this study are available from the corresponding author upon request. Source data are provided with this paper
IX Code availability
Key algorithms of theoretical calculations are fully clarified in the Supplementary Information.

### v2

@@ Abstract
A further characteristic of this system is the emergence of an interlayer excitonic phase at $\nu_{\rm tot}=-1$. In this regime the top sheet forms a conventional integer quantum Hall state at $\nu_{\rm top}=-2$, while the middle and bottom sheets are simultaneously half filled, creating conditions for excitonic coherence across adjacent layers. The resulting suppression of $R_{xx}$ resembles excitonic behavior reported previously in twisted bilayer and double bilayer graphene, but here it coexists with a fully filled outer layer, providing a trilayer realization of this phenomenon. The magnetic-field dependence of the feature is consistent with an interaction-driven origin. In summary, large-angle TTG combines conventional quantum Hall features with interlayer correlated phases, illustrating how trilayer systems with tunable layer asymmetry can host spin-resolved edge transport and excitonic order.
I Supporting Information
Methods, Additional Experimental Data, Band structure calculations, Hartree-Fock calculations, Non-local measurement
II Acknowledgement
The work from DGIST was supported by the National Research Foundation of Korea (NRF) (Grant No. RS-2025-00557717, RS-2023-00269616) and the Nano and Material Technology Development Program through the National Research Foundation of Korea (NRF) funded by Ministry of Science and ICT (No. RS-2024-00444725). We also acknowledge the partner group program of the Max Planck Society. Part of this work was supported by Global Partnership Program of Leading Universities in Quantum Science and Technology (RS-2025-02317602). G. Y. C. is supported by Samsung Science and Technology Foundation under Project Number SSTF-BA2401-03, the NRF of Korea (Grants No. RS-2023-00208291, RS-2024-00410027, 2023M3K5A1094810, RS-2023-NR119931, RS-2024-00444725, RS-2023-00256050, IRS-2025-25453111, RS-2025-08542968) funded by the Korean Government (MSIT), the Air Force Office of Scientific Research under Award No. FA23862514026, and Institute of Basic Science under project code IBS-R014-D1. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) , the CREST (JPMJCR24A5), JST and World Premier International Research Center Initiative (WPI), MEXT, Japan. This work at UOS was supported by the National Research Foundation of Korea (NRF) through Grant No. RS-2023-00249414 (N.L.) and Grant No. NRF-2020R1A5A1016518 (J.J.). Computational resources were provided by KISTI under Grant No. KSC-2022-CRE-0514 and by the Urban Big Data and AI Institute (UBAI) at UOS.